A package of 6 fuses are tested where the probability an individual fuse is defective is 0.5
What is the probability that one fuse will be defective
1
Expert's answer
2020-05-18T18:27:50-0400

Binomial distribution problem with "P=0.05" and "n=6"

"P(=1) =0.05 \times{(1-0.05)^5} \times {6}=0.2321342"